# Digital Twin-Empowered Task Assignment in Aerial MEC Network: A Resource Coalition Cooperation Approach with Generative Model

### v3

@@ III-B DT Model
In this paper, the communication between MED and UAV (M2U) is designed to obtain the MED task parameters to build a DT model. In addition, the communication between UAV and airship (U2A) is designed to transmit the DT modeling parameters of MED and UAV. The above communication process all adopts the communication method of time division multiple access. To simplify the analysis, this paper defines the system model as a synchronous periodic time slot system, and the duration of each time slot $t$ is $\tau$. Let the MED set be $M=\left\{{1,2,\cdots,i,\cdots,m}\right\}$. $m$ means the maximum sequence number of MED. The $i$th MED is represented as $MED_{i}$. Then the DT of $i$th MED in time slot $t$ is
@@
The set of UAV is denoted as $V=\left\{{1,2,\cdots,j,\cdots,n}\right\}$. $n$ means the maximum sequence number of UAV. Let $UAV_{i}$ be the $i$th UAV, DT of $UAV_{j}$ within time slot $t$ is given as follows
@@
$\vartheta^{t}_{i}$ Computational complexity
@@ IV-A Definition of Task Assignment
Peer-to-peer pooling: All the nodes equipped with various resources in the aerial MEC network are regarded as equal, which builds a resource pool among the nodes. Furthermore, the resources of the nodes can be used through cooperative games, and their resource capabilities will be uniformly scheduled.
Low-to-high scheduling: The nodes in the aerial MEC network are composed of end users (i.e., MEDs), edge servers (i.e., UAVs), and cloud servers (i.e., airship). The tasks in our proposed network can be executed in low-to-high manner according to task request: 1) one-to-one executing; 2) one-to-many assigning. In this paper, we focus on scheduling resources on multiple nodes to complete the tasks.
One-to-many assigning: In existing task offloading strategies, the tasks are usually offloaded to a powerful edge server or a remote cloud server for increasing computing efficiency (i.e., one-to-one executing). However, in the proposed aerial MEC network, the tasks can be assigned to many edge servers.
@@ IV-D Coalition Participant Utility
where $f^{t}_{j,max}$ is maximum available computing resources of $UAV_{j}$. $C_{6}$ indicates that the $UAV_{j}$ can decide whether to accept the task assignment, and if it accepts, the UAV cannot be overclocked when executing its task. $C_{7}$ denotes that the time it takes for the coalition to complete the accepted task must meet the latency requirements of the $MED_{i}$.
In our proposed method, the UAV acts as an aerial base station for providing computing services. Therefore, the utilization of the computing resources of the UAV must be considered when obtaining the maximum participant utility. The utilization of the computing resource $mu$ by each participant in the coalition is designed as follows
@@ V Resource Coalition Cooperation Approach based on DT
In this section, we first introduce a definition of coalition formation, including Pareto optimal, coalition combining, and coalition dividing rule. Next, we formulate the resource coalition cooperation algorithm based on DT.
@@ V-B Resource Coalition Cooperation Algorithm
  19: According to the Step 4, coordinate the implementation of the resource coalition cooperation. The user then assigns the disassembled tasks to multiple UAVs for computing.
@@
Figure 4: DT-empowered TU coalition game method.
@@ VI Performance Analysis
In this section, the stability and convergence of TU coalition game are illustrated. Then, we define the fidelity of DT model and analyse its deviation.
VI-A Stable Coalition Definition and Proof
@@ VI-C DT Deviation Analysis
High-fidelity DT model relies on real-time interactions between physical space and virtual space. The computing resources of the UAV change dynamically, so that its computing resource $f_{j}$ cannot be accurately obtained. Therefore, the actual computing resource of $UAV_{j}$ is ${\tilde{f}}_{j}=f_{j}+\mathrm{\Delta}f_{j}$, where $\mathrm{\Delta}f_{j}$ is the deviations. In order to measure the accuracy of DT model reflecting physical object parameters, we define the fidelity ${{\mathcal{F}}_{\mathcal{N}}}=1-\left|\Delta\mathcal{N}/\mathcal{\tilde{N}}\right|$ as the deviation between the estimated value in the DT model and the true value of the corresponding physical object status quantity [44]. That is, the smaller the deviation between the DT model parameters and the physical real values, the higher the fidelity of the DT model generated in the formal update phase. ${{\mathcal{F}}_{{{f}_{j}}}}=1-\left|\Delta{{f}_{j}}/{{\tilde{f}}_{j}}\right|$ is the fidelity of computing resource.
@@ VII-D Resource Utilization Results and Analysis
The utilization computing resources for different numbers of UAVs are given in Fig. 9. It can be seen that the coalition game approach proposed in this paper always outperforms the grand coalition and Nash equilibrium approaches. The reason for this is that the coalition game approach requires only some UAVs to act as coalition participants to provide services for task assignment, and the computing resource provided by each UAV are iteratively optimized. In contrast, the grand coalition approach is to provide all UAVs with services for task offloading, and the resources allocated to each UAV are time-invariant, and the various onboard resources cannot be dynamically adapted to the changes of the task. At the same time, as the number of UAVs participating in the task grows, the resource utilization of the grand coalition approach is inferior to that of the proposed approach. Unlike the previous two cooperative game methods, the Nash equilibrium focuses on maximizing the benefits of each participant’s coalition and does not form resource collaboration between them. Therefore, although the multi-UAV resource utilization based on the Nash equilibrium improves compared to the grand coalition, it lacks an effective resource cooperation mechanism between multi-UAVs, and with the incremental increase in the number of UAVs, the resource utilization based on the method does not improve significantly.
@@ VII-E Energy Consumption Results and Analysis
The results of the total energy consumption of all UAVs for different computational complexities are given in Fig. 10. From this figure, it can be observed that the total energy consumption of all UAVs gradually increases as the threshold value of the task execution latency increases. This is because as the computational complexity increases, the UAV is required to invest more arithmetic power in processing and requires more computation latency, which in turn leads to the computing energy consumption being proportional to the computational complexity, and obviously, the total energy consumption is also elevated. At low computational complexity ($\tau$ = 50), given that the coalitional gaming approach requires only partial UAV participation to accomplish this type of task, the total energy consumption has a significant advantage over the remaining two types of approaches. However, as the computational complexity of the task increases ($\tau$ = 300), more UAVs are required to participate in the task execution, which leads to a decrease in the energy performance gap between these three types of methods. The coalition game and the Nash equilibrium-based methods are based on the premise of improving the utility of the UAVs, and thus the energy performance of the two approaches converges to the same level.
@@ VIII Conclusion
In this paper, we propose a DT-empowered task assignment based on resource coalition cooperation with a GM in an aerial MEC network that realizes the mapping of physical space resources to a resource pool in virtual space. Furthermore, to improve the energy consumption of each coalition participant as little as possible and further adapt to the MED task assignment request to maximize the utility of the coalition participants, our proposed method optimizes the bandwidth and computing resource cooperation between UAVs by establishing a coalition game model with a GM so as to realize the sharing and optimal deployment of different resources. The advantages of a coalition game with a GM can be leveraged to improve iterative efficiency in decision-making. The simulation results show that, compared with benchmark algorithms, the proposed method increases resource utilization and the average utility of the coalition participants, which not only realizes the dynamic matching of resource reserves of an aerial MEC network with demand changes but also reduces energy consumption. The proposed method can jointly optimize the satisfaction of UAVs and the interactions of U2U.

### v4

@@ III-B DT Model
In this paper, the M2U is also designed to obtain the MED task parameters to construct a DT model. In addition, the communication between UAV and airship (U2A) is designed to transmit the DT modeling parameters of MED and UAV. The above communication process all adopts the communication method of time division multiple access. To simplify the analysis, this paper defines the system model as a synchronous periodic time slot system, and the duration of each time slot $t$ is $\tau$. Let the MED set be $M=\left\{{1,2,\cdots,i,\cdots,m}\right\}$. $m$ means the maximum sequence number of MED. The $i$th MED is represented as $MED_{i}$. Then the DT of $i$th MED in time slot $t$ is
@@
The set of UAV is denoted as $V=\left\{{1,2,\cdots,j,\cdots,n}\right\}$. $n$ means the maximum sequence number of UAV. Let $UAV_{j}$ be the $j$th UAV, DT of $UAV_{j}$ within time slot $t$ is given as follows
@@
$\vartheta$ Computational complexity
@@ IV-A Definition of Task Assignment
Peer-to-peer pooling: All UAVs equipped with various resources in the aerial MEC network are considered equal, contributing to the establishment of a resource pool in the airship. Furthermore, the resources of the UAVs can be used through cooperative games, and their resource capabilities will be uniformly scheduled.
Low-to-high scheduling: The nodes in the aerial MEC network are composed of end users (i.e., MEDs), edge servers (i.e., UAVs), and cloud servers (i.e., airship). The tasks in our proposed network can be executed in low-to-high manner according to task request: 1) one-to-one executing; 2) one-to-many assigning. In this paper, we focus on scheduling resources on multiple UAVs to complete the tasks.
One-to-many assigning: In existing task offloading strategies, the tasks are usually offloaded to a powerful edge server or a remote cloud server for increasing computing efficiency (e.g., One-to-one task offloading, see Fig. 1). However, in the proposed aerial MEC network, the tasks can be assigned to many edge servers.
@@ IV-D Coalition Participant Utility
where $f^{t}_{j,max}$ is maximum available computing resources of $UAV_{j}$. $C_{6}$ indicates that the $UAV_{j}$ can decide whether to accept the task assignment, and if it accepts, the UAV cannot be overclocked when executing its task. $C_{7}$ denotes that the time it takes for the coalition to complete the accepted task must meet the latency constraints of the $MED_{i}$.
In our proposed method, the UAV acts as an aerial base station for providing computing services. Therefore, the utilization of the computing resources of the UAV must be considered when obtaining the maximum participant utility. The utilization of the computing resource $\mu$ by each participant in the coalition is designed as follows
@@ V Resource Coalition Cooperation Approach based on DT
In this section, we first introduce a definition of coalition formation, including Pareto optimal, coalition combining, and coalition dividing. Next, we formulate the resource coalition cooperation algorithm based on DT.
@@ V-B Resource Coalition Cooperation Algorithm
  19: According to the Step 4, coordinate the implementation of the resource coalition cooperation. The user then assigns the tasks to multiple UAVs for computing.
@@
Figure 4: DT-empowered resource coalition cooperation.
@@ VI Performance Analysis
In this section, the stability and convergence of coalition are illustrated. Then, we define the fidelity of DT model and analyse its deviation.
VI-A Stability Definition and Proof
@@ VI-C DT Deviation Analysis
High-fidelity DT model relies on real-time interactions between physical space and virtual space. The computing resources of the UAV change dynamically, so that its computing resource $f_{j}$ cannot be accurately obtained. Therefore, the actual computing resource of $UAV_{j}$ is ${\tilde{f}}_{j}=f_{j}+\mathrm{\Delta}f_{j}$, where $\mathrm{\Delta}f_{j}$ is the deviations. In order to measure the fidelity of DT model reflecting physical object parameters, we define the fidelity ${{\mathcal{F}}_{\mathcal{N}}}=1-\left|\Delta\mathcal{N}/\mathcal{\tilde{N}}\right|$ as the deviation between the estimated value in the DT model and the true value of the corresponding physical object status quantity [44]. That is, the smaller the deviation between the DT model parameters and the physical real values, the higher the fidelity of the DT model generated in the formal update phase. ${{\mathcal{F}}_{{{f}_{j}}}}=1-\left|\Delta{{f}_{j}}/{{\tilde{f}}_{j}}\right|$ is the fidelity of computing resource.
@@ VII-D Resource Utilization Results and Analysis
The utilization of computing resources for different numbers of UAVs are given in Fig. 9. It can be seen that the coalition game approach proposed in this paper always outperforms the grand coalition and Nash equilibrium approaches. The reason for this is that the coalition game approach requires only some UAVs to act as coalition participants in providing services for task assignment, while the computing resource cooperation strategies are iteratively optimized. In contrast, the grand coalition approach is to provide all UAVs with services for task offloading, and the resources allocated to each UAV are time-invariant, and the various onboard resources cannot be dynamically adapted to the changes of the task. At the same time, as the number of UAVs participating in the task grows, the resource utilization of the grand coalition approach is inferior to that of the proposed approach. Unlike the previous two cooperative game methods, the Nash equilibrium focuses on maximizing the benefits of each participant’s coalition and does not form resource collaboration between them. Therefore, although the multi-UAV resource utilization based on the Nash equilibrium improves compared to the grand coalition, it lacks an effective resource cooperation mechanism between multi-UAVs, and with the incremental increase in the number of UAVs, the resource utilization based on the method does not improve significantly.
@@ VII-E Energy Consumption Results and Analysis
The results of the total energy consumption of all UAVs for different computational complexities are given in Fig. 10. From this figure, it can be observed that the total energy consumption of all UAVs gradually increases as the threshold value of the task execution latency increases. This is because as the computational complexity increases, the UAV is required to invest more arithmetic power in processing and requires more computation latency, which in turn leads to the computing energy consumption being proportional to the computational complexity, and obviously, the total energy consumption is also elevated. At low computational complexity ($\vartheta$ = 50), given that the coalitional gaming approach requires only partial UAV participation to accomplish this type of task, the total energy consumption has a significant advantage over the remaining two types of approaches. However, as the computational complexity of the task increases ($\vartheta$ = 300), more UAVs are required to participate in the task execution, which leads to a decrease in the energy performance gap between these three types of methods. The coalition game and the Nash equilibrium-based methods are based on the premise of improving the utility of the UAVs, and thus the energy performance of the two approaches converges to the same level.
@@ VIII Conclusion
In this paper, we propose a DT-empowered task assignment based on resource coalition cooperation with a GM in an aerial MEC network that realizes the mapping of physical space resources to a resource pool in virtual space. Furthermore, to improve the energy consumption of each coalition participant as little as possible and further adapt to the MED task assignment request to maximize the utility of the coalition participants, our proposed method optimizes the bandwidth and computing resource cooperation between UAVs by establishing a coalition game model with a GM so as to realize the sharing and optimal deployment of different resources. The advantages of a coalition game with a GM can be leveraged to improve iterative efficiency in decision-making. The simulation results show that, compared with benchmark algorithms, the proposed method increases resource utilization and the average utility of the coalition participants, which not only realizes the dynamic matching of resource reserves of an aerial MEC network with demand changes but also reduces energy consumption. The proposed method can jointly optimize the utility of UAVs and the interactions of U2U.